Consider the following information showing the preference among boys and girls for orange juice and apple juice.

Boys and Apple Juice 10
Boys and Orange Juice 20
Girls and Apple Juice 22
Girls and Orange Juice 18

If the computed Chi Square statistic is 3.24, obtain the p-value.
Answer #1

Ans:

Observed(fo)
Apple orange Total
Boys 10 20 30
Girls 22 18 40
Total 32 38 70
Expected(fe)
Apple orange Total
Boys 13.71 16.29 30
Girls 18.29 21.71 40
Total 32 38 70
Chi square=(fo-fe)^2/fe
Apple orange Total
Boys 1.01 0.85 1.85
Girls 0.75 0.64 1.39
Total 1.76 1.48 3.24

Test statistic:

Chi square=3.24

df=(2-1)*(2-1)=1

p-value=CHIDIST(3.24,1)=0.0719